OpenStudy (anonymous):

Simply the expression. -9+7z+2-3z=

OpenStudy (mathstudent55):

Combine like terms together. Terms that have z in them are like terms. Terms that don't have z are like terms. -9 + 7z + 2 - 3z = -9 + 2 + 7z - 3z =

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

$$a+b=b+a$$

OpenStudy (mathstudent55):

What is -9 + 2?

OpenStudy (anonymous):

-11

OpenStudy (mathstudent55):

That is -9 - 2. What is -9 + 2?

OpenStudy (anonymous):

+11

OpenStudy (mathstudent55):

No. Let's say the temperature was -9 degrees and it went up 2 degrees. What is the temperature now?

OpenStudy (anonymous):

9

OpenStudy (mathstudent55):

@SeanAcedia What does 7 + (-2) have to do with this other than adding more confusion?

OpenStudy (anonymous):

ok,iam confused

OpenStudy (mathstudent55):

@sjackson Let's say the temperature was -9 degrees and it went up 2 degrees. What is the temperature now?

OpenStudy (anonymous):

11 correct?

OpenStudy (mathstudent55):

No. The temperature was not 9 degrees, It was -9 degrees. Then it went up 2 degrees. What is it now?

OpenStudy (anonymous):

-11

OpenStudy (mathstudent55):

Here's another way of thinking of negative numbers. I owe you 9 dollars, so you can think of me having -9 dollars. Now I pay you back 2 dollars. How much do I owe you now?

OpenStudy (anonymous):

-7

OpenStudy (mathstudent55):

Right. -9 + 2 = -7 Now the part with the variable z: What is 7z - 3z = ?

OpenStudy (anonymous):

4z

OpenStudy (mathstudent55):

Right. So the final answer is: -9 + 7z + 2 - 3z = -7 + 4z

OpenStudy (anonymous):

11z

OpenStudy (mathstudent55):

Remember what I wrote at the top. You can only combine like terms. Like terms have the same variable part. Terms with z are like terms. Terms without z are like terms. Terms with z and terms without z are not like terms and can't be combined together. The final answer is -7 + 4z You can't do anything else to this.

OpenStudy (anonymous):

there not like terms if a number has a letter in it and another number doesnt have a letter in it they cant be put together.

OpenStudy (skullpatrol):

Why 11z?

OpenStudy (mathstudent55):

@skullpatrol It's the incorrect addition of -7 and 4

OpenStudy (skullpatrol):

@sjackson -7 + 4z can not be simplified further, do you know why?